# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v8

@@ 1 Introduction
We note that there are offensive words in this paper, including in figures and tables, and especially in sections discussing extant tokens, bias, and alignment.
@@ 1.2 The Distributional Hypothesis
Because input is flattened into 1D sequences of tokens, the LLM has to reconstruct any syntax syntagmatically. Syntax can expose word boundaries through recombination. In fact, that many successful models use vocabularies that incorporate a leading or trailing space with relevant tokens supports this interpretation. The whitespace as its own token would change the distribution, whereas incorporating whitespace with words means the sentence is segmented as a person would: with word boundaries. Although not well understood, people “can segment words and other constituents from each other with astonishing precision”, through mechanisms such as the lengthening of word-initial consonants Blum et al. (2024). That the space is often attached to the corresponding token by tokenization algorithms as well is suggestive of emergent harmony between the tokenizer and what people seem to be doing, something we return to in Section 2.1.252525We suspect this emergent harmony comes from the shared aspects of a human-honed language implementation; looking for reasonably optimized engineering solutions also naively pushes the computer science and software engineering relevant to LLMs towards linguistic theory, as in Stafford Beer’s famous phrase, “the purpose of a system is what it does”.
@@ 2.1 Exemplar tokenizations
A tokenization algorithm uses the training data it is given to determine what pieces (tokens) will be needed for future output. The popular tokenization design choice of maximizing orthographic efficiency allows as much data as possible to be processed by the LLM, presumably the motivation behind this strategy’s ubiquity. We trained one such tokenizer, a Byte-Pair Encoding (BPE) tokenizer Anthony Moi and Nicolas Patry (2023), with varying vocabulary sizes and training data to explore the details of the process, revealing some alignment between linguistic features and information processing by showing the human-like stages the tokens pass through. As a caveat, the training data chosen also determines the tokens; once vocabulary size becomes relatively large compared to the training data, tokens become too specific. The BPE implementation we used here was character-level, not byte-level, and we retained some (five) LLM-related special tokens as default items in the vocabulary. As a result, whatever our vocabulary size was fixed to, $N$, was in reality N5N-5, because it already included a few irrelevant (to us) tokens. We did not explicitly call any pre-tokenizers, in particular, the Whitespace pre-tokenizer, because that would mean the tokenizer would have automatically split on whitespace rather than treating the text agnostically as a continuous string of characters HuggingFace (2025). BPE is by no means the only tokenization algorithm used in LLMs, as aforementioned (Section 1.1), but our goal is to understand what kind of consequences follow for LLMs given that we have made tokenization – of any kind – an integral part of their architecture. Especially of interest to us are the implications of popular tokenization approaches that focus on parsing vast corpora of data efficiently, such as BPE, rather than on another goal, such as mimicking human language (which might make sense given the human-brain-inspiration of artificial neural nets and their application in LLMs to human language as primary data source) or on strategically designing cognition-scaffolding semantic primitives. [^28] The motivation underlying the creation of the Toki Pona conlang, and which would seem well-suited to neurosymbolic approaches.,  [^29] Future work could compare systematically across tokenization strategies.
@@
Furthermore, we can see that words are not the only linguistic unit that the tokenization algorithm can surface, meaning those units, too, are informatively optimal under specific conditions (those in place when they are surfaced by the tokenization algorithm). We see a variety of linguistic structures including morphemes (free and bound, productive affixes), words, phrases, and sentences (or utterances). See Figures 2627: the X label means unknown/ couldn’t label, so seeing that move down in the distribution of category counts reflects that entities not matching those sought by the parsing package (human-meaningful syntactic units e.g., phrases such as noun phrases (NP)) are becoming less likely—in other words, that tuning the vocabulary size can cause the tokenizer to surface more human-like syntactic structures (as well as words).
@@
That we see human-meaningful linguistic units appear as tokens evinces a relationship between informative efficiency (BPE is a compression algorithm, so entropy-reducing Sennrich, Haddow, and Birch (2016); Gage (1994)) and the structures in human language, in line with a growing body of work on how such forces shape language Gibson et al. (2019). As Futrell and Hahn, 2024 put it, “the structure of human language may have evolved to minimize cognitive load while maximizing communicative expressiveness” Futrell and Hahn (2024) – echoing the motivation of maximizing orthographic efficiency in order to pack as much text as possible into the LLM. [^33] Caveat: we are moving fluidly between talking about language and text. Of course, orthography does not equal speech, but it is not unrelated.
@@
We note that the pressures of orthographic efficiency—akin to popular tokenization algorithms—are reflected in world writing systems through the development of ligature, and in the development of contracted/ Grade 2 braille. More broadly, we can see forces of efficiency at work in language at all levels in the general principle lectio difficilior potior; for example, a non-trivial driver of language change over time is that people migrate a language towards sounds that are physically easier to make Cowen-Breen et al. (2023).
If our interpretation of what is happening with the DH is right, then models should perform better at the human-like vocabulary stages of tokenization (see Figure 3), at least through increasing the salience of useful distributions, and potentially also by providing suitable receptacles as semantic primitives for information derived from human cognition. The importance of words is suggested by the parallel word-boundary marking observed in both people and LLMs, and hints at a definition: perhaps, if a word is the unit of language best suited to our ability to learn distributionally from the linguistic signal, then models should perform best when provided easy access to plentiful word tokens.
@@ 2.3 Extispicy: exemplar tokens
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^53] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions545454e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Figure 41).
@@ 3.3 Flattening
The tokens form the semantic primitives for the model. They likely need to be well-formed receptacles for the kind of semantic information that makes up human world models, as that is a primary role of words for us.808080“A primary function of the lexicon is to encode novel concepts and ideas as they emerge over time” Xu et al. (2024). The Sapir-Whorf hypothesis in general, nameability (Lupyan’s Label-feedback hypothesis), and even theories from language acquisition like fast-mapping offer explanations for what kind of processes might need such structures Scholz et al. (2024); Lupyan (2012); Gelman and Brandone (2010). [^81] Although LLMs have something like MVP language, how that maps to the intersubjective or intrasubjective competence humans derive from language is unknown.
@@ 3.5 Bias, alignment, and ethics
Given the previous points, it is worth flagging that although they make up small portions of the total vocabulary, it is not uncommon for bad words to appear as tokens (especially since many new models remix or fine-tune existing models, and often incorporate an existing vocabulary off-the-shelf, without close inspection). [^94] LLMs are “so expensive to retrain from scratch that they are typically initialized with pre-trained models” Shumailov et al. (2024). In addition, changing the vocabulary would typically mean throwing away the initial training that had been done Bostrom and Durrett (2020). For example, 126 (out of 965) “bad words” are tokens in GPT-4o. We found some tokens that seem particularly likely to be vehicles for information that would work counter to alignment goals; for example, “cunt”, “slut”, “bitch”, “negro”, and “coon” are GPT-4o tokens, as are “iggers”, “igger”, and “IGGER”. The last three are representative of a class of tokens under suspicion as components of unsavory words, which, as aforementioned, could still carry that same content. [^95] In more detail: consider “igger”, “iggers”, and “IGGER” in GPT-4’s vocabulary (cl100k_base). Of course, those can be subwords of multiple words. For example, “bigger” and many variations of “trigger” are tokens, while “digger” is not. If “igger”, etc., is a useful subword frequently seen in other words, it would be included as a token. Depending on the details of the tokenization algorithm, inclusion of words containing it may or may not continue to contribute to its presence in the final vocabulary. That is, “igger” would be added to the vocabulary at some point, then “trigger” and “bigger”, and then, depending on the algorithm, a process may or may not remove “igger” if its value has decreased. However, there are a relatively small number of words containing it as a subword—some of those words are variants of the n-word, and several of the other, more mundane words are present as their own tokens. The presence of the subword “IGGER” in an entirely capitalized form is particularly suspicious (of low-quality or problematic content [^96] Potentially up to and including the repulsive dregs of our collective written works - racism, hatred, bigotry, etc. related to the n-word or to “triggered”) given orthographic conventions on the internet. [^97] Meaning both emphatic capitalization and the use of capitalization in memes. (See Table 11, Figures 3738.) The suggested content here is compatible, in our view, with features found in Templeton, et al., 2024: their analysis of the Claude 3 Sonnet transformer model at the feature level—above the token level—turned up concerning categories we could imagine being amongst just the sort of data we’re worried about making its way from the internet into the model, including “Phrases indicating profanity, vulgarity, obscenity or offensive language”; “Racist slurs and offensive language targeting ethnic/racial groups, particularly the N-word”; “Derogatory slurs, especially those targeting sexual orientation and gender identity”  Templeton et al. (2024). [^98] And “Offensive, insulting or derogatory language, especially against minority groups and religions”; “Racist claims about crime”; “Mentions of violence, malice, extremism, hatred, threats, and explicit negative acts”. They also found information on famous people (which goes to a discussion of proper nouns in this paper and celebrities and Twitter in the in-progress Cramer, et al. Cramer et al. (2025)). They report that the features they found respond to and cause abstract behaviors Templeton et al. (2024). We think it is plausible that these features are, at least (and probably at most) in part, due to the mechanisms we’ve described.
@@
These same issues (bias, harmful content) can reasonably be assumed to persist throughout the model, since they are also attested in output, even after alignment. Hofmann et al., 2024 found that “HF [human feedback] training obscures the racism on the surface, but the racial stereotypes remain unaffected on a deeper level” Hofmann, Kalluri, and Jurafsky (2024). It is worth noting that many current alignment strategies would not be expected to fully redress the pathways for unwanted content that we’ve described. [^99] Current prominent alignment strategies such as RLHF (Reinforcement Learning from Human Feedback) are unlikely to address the initial embedding matrix, although they can impact the content of the token vectors as they move through the model, since they are aimed at adjusting the model’s weights Stiennon et al. (2022). Even assuming alignment went exactly as intended and was as successful as possible, unwanted information could plausibly find a backdoor into the model through the initial token vectors during early training/ tokenization and remain untouched. A similar argument is put forward in Lee et al., 2025 with respect to the inadequacy of current unlearning techniques: “Even when achieving perfect behavioral agreement with an unlearning oracle, models retain latent capabilities that can be easily restored through finetuning” Lee et al. (2025).
@@
Beyond alignment, based on the the in-domain problem and the ouroboros: archaic, obscure, and obsolete words are likely not going to be tokenized.100100100Our results are consistent with this finding from Zhemchuzhina et al., 2022 Zhemchuzhina, Filippov, and Yamshchikov (2022)! How is that going to contribute to loss of beautiful cultural artifacts? Beyond buildings and even texts, language itself is a heritage site.101101101For example, peruse the Phrontistery Chrisomalis (1996-2021)! Similar arguments have been made at the narrative and cultural level, beyond the individual word or phrase: Generative AI acts as a “crucible of mediocrity” Ruiz, Zimmerman, and team (2023), as a regression to the mean Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).102102102We speculate that the probability mass of the logits corresponding to the subset of tokens involved in inference is likely to be relatively high, and that this is related to why models sometimes behaviourally collapse into inane repetition. More broadly, that repetition is a symptom of the iterative concentration of future plausibility based on previous plausibility, which is effectively a one-way ratchet Zimmerman et al. (2024). Imagine a hypothetical LLM generating output based on every previous utterance, including what it, itself, has said. As an analogy, compare sequence A005150 Foundation Inc. (2024a) to sequence A363054 Foundation Inc. (2024b): the initial seeds are obviously different, but as the output length grows, those differences are essentially lost to time (history being deeper on the right side of each term), and the sequences feel increasingly indistinguishable, locked in to reiterating the same sorts of patterns. It tends towards being normative and generic, which is undercut by creativity, novelty, and diversity Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).103103103“Beyond achieving linguistic coherence, beyond resolving the ‘hallucinations’ and errors, it is the performance of the generic and the normative, sounding ‘right,’ that sells AI tools… As currently designed, generative AI tend towards being generic, centrist, normative, and banal. Hallucinations undermine that, but so do cultural and narrative diversity” Gillespie (2024). This is, of course, partly a consequence of the fuzzing up of grammaticality or frequency with truth Zimmerman et al. (2024)!104104104McCoy et al., 2024, provides support as to (1) our characterization of plausible text generation as intertwining frequency and truth and (2) the theoretical importance of being able to modulate the primacy of the plausible text generation objective function, and the limitations of our current attempts in that vein NTP vs “reasoning”, efficacy of attempts to modulate primacy of NTP: “Specifically, o1 – like previous LLMs – is sensitive to the probability of examples and tasks, performing better and requiring fewer ‘thinking tokens’ in high-probability settings than in low-probability ones. These results show that optimizing a language model for reasoning can mitigate but might not fully overcome the language model’s probability sensitivity” McCoy et al. (2024b).,105105105For more discussion of the DH, and regime 1 vs. regime 2 learning (pattern-matching vs. learning from abstractions; hearing a novel word in context vs. reading a dictionary definition), see  Zimmerman (2025).
@@
Current LLMs have trouble with unusual data, at all scales.106106106Language, too, has this tendency inalienably within it, via its fungibility. Information technologies homogenize: that is both the benefit and price of fungibility. Fungibility is inherently degrounding, and degrounding is inherently homogenizing; it entails a loss of information or meaning, but a gain in utility. See first author’s thesis for more discussion of the construction of meaning along the fungibility-grounding spectrum Zimmerman (2025). Unusual data, data in the tails of the distribution, is important, both for cultural reasons and practical reasons of model performance and collapse. This is especially worth considering in the light of proposals to use LLMs to generate synthetic data (the ouroborous).107107107Shumailov et al., 2024 found that “in learning tasks in which the tails of the underlying distribution matter, one needs access to real human-produced data” as opposed to relying on LLM-generated data. They found that “access to the original data distribution is crucial” in order to prevent model collapse in subsequent generations Shumailov et al. (2024). Then, an initial tokenization step that curtails a model’s access to the tails of the attested word use distribution by omitting longer and rare words could also contribute to model collapse or other performance issues. Tokenization is a deterministic (though context-dependent) process, from within the perspective of a single algorithm being applied to a single fixed corpus. But tokenization as a cultural artifact in a larger system is less transparent.108108108Research continues to emerge that provides additional support for this argument Yakura et al. (2025).

### v9

@@ 1 Introduction
Please note that there are offensive words in this paper, including in figures and tables, and especially in sections discussing extant tokens, bias, and alignment.
@@ 1.2 The Distributional Hypothesis
Because input is flattened into 1D sequences of tokens, the LLM must reconstruct syntactic structure syntagmatically. A primary cue for this reconstruction is the word boundary, which syntax exposes through the recombination of discrete units. Incorporating whitespace directly into word tokens provides the model with a segmentation that better aligns with human intuition as to discrete words; people “can segment words and other constituents from each other with astonishing precision” through mechanisms such as the lengthening of word-initial consonants Blum et al. (2024). Many successful models use vocabularies that incorporate a leading or trailing space with relevant tokens, suggesting an emergent harmony between the tokenization process and what people seem to be doing, something we return to in Section 2.1.252525Mandelbrot argued that word-like segmentation is an optimal coding strategy when trying to minimize transmission cost in a toy model of language Mandelbrot (1953, 1954), a principle now algorithmically realized by BPE and routinely surfaced in the vocabularies of existing, performant LLMs. We suspect this emergent harmony comes from the shared aspects of a human-honed language implementation, notably including the principle of least effort, which we return to in Section 2.1; looking for reasonably optimized engineering solutions also pushes the computer science and software engineering relevant to LLMs towards linguistic theory, as in Stafford Beer’s famous phrase, “the purpose of a system is what it does”.
@@ 2.1 Exemplar tokenizations
A tokenization algorithm uses the training data it is given to determine what pieces (tokens) will be needed for future output. The popular tokenization design choice of maximizing orthographic efficiency allows as much data as possible to be processed by the LLM, presumably the motivation behind this strategy’s ubiquity. We trained one such tokenizer, a Byte-Pair Encoding (BPE) tokenizer Anthony Moi and Nicolas Patry (2023), with varying vocabulary sizes and training data to explore the details of the process, revealing some alignment between linguistic features and information processing by showing the human-like stages the tokens pass through. As a caveat, the training data chosen also determines the tokens; once vocabulary size becomes relatively large compared to the training data, tokens become too specific. The BPE implementation we used here was character-level, not byte-level, and we retained some (five) LLM-related special tokens as default items in the vocabulary. As a result, whatever our vocabulary size was fixed to, $N$, was in reality N5N-5, because it already included a few irrelevant (to us) tokens. We did not explicitly call any pre-tokenizers, in particular, the Whitespace pre-tokenizer, because that would mean the tokenizer would have automatically split on whitespace rather than treating the text agnostically as a continuous string of characters HuggingFace (2025). BPE is by no means the only tokenization algorithm used in LLMs, as aforementioned (Section 1.1), but our goal is to understand what kind of consequences follow for LLMs given that we have made tokenization – of any kind – an integral part of their architecture. Especially of interest to us are the implications of popular tokenization approaches that emphasize the processing of vast corpora of data efficiently, such as BPE, rather than another goal, such as replicating human language (which might make sense given the human-brain-inspiration of artificial neural nets and their application in LLMs to human language as primary data source) or on strategically designing cognition-scaffolding semantic primitives. [^28] The motivation underlying the creation of the Toki Pona conlang, and which would seem well-suited to neurosymbolic approaches.,  [^29] Future work could compare systematically across tokenization strategies.
@@
Furthermore, we can see that words are not the only linguistic unit that the tokenization algorithm can surface, meaning those units, too, are informatively optimal under specific conditions (those in place when they are surfaced by the tokenization algorithm). We see a variety of linguistic structures including morphemes (free and bound, productive affixes), words, phrases, and sentences (or utterances). See Figures 2627: the X label means unknown/ couldn’t label, so seeing that move down in the distribution of category counts reflects that entities not matching those sought by the parsing package (human-meaningful syntactic units e.g., phrases such as noun phrases (NP)) are becoming less likely—in other words, that tuning the vocabulary size can cause the tokenizer to surface more human-like syntactic structures (as well as words), across linguistic scales.
@@
The emergence of human-meaningful linguistic units as tokens in LLMs evinces a relationship between informative efficiency (BPE is a compression algorithm, thus entropy-reducing Sennrich, Haddow, and Birch (2016); Gage (1994)) and the structures of human language, consistent with prior work on efficiency pressures in language Gibson et al. (2019). As Futrell and Hahn, 2024 note, “the structure of human language may have evolved to minimize cognitive load while maximizing communicative expressiveness” Futrell and Hahn (2024) – a principle echoed in the design of LLM tokenization, which often prioritizes a compact representation of the training text. [^33] Caveat: we are moving fluidly between language and text; orthography is not speech, but it is related.
@@
Orthographic efficiency pressures – mirrored in tokenization algorithms – also appear in world writing systems through ligatures and contracted/Grade 2 braille. More generally, we can see forces of efficiency at work in language at all levels in the principle lectio difficilior potior, or the principle of least effort (PLE); [^34] Lectio difficilior potior (LDP) can be seen as a special case of PLE in language transmission. If the connection between LDP and PLE has not been explicitly formalized, then doing so through computational modeling or other analysis could be useful future work. for example, a non-trivial driver of language change over time is that people migrate a language towards sounds that are physically easier to make Cowen-Breen et al. (2023). Observing these efficiency pressures re-emerge in the context of LLMs provides a striking demonstration that optimization principles long hypothesized for human language can also shape its implementation in artificial systems.
If our interpretation of what is happening with the DH is correct, then models should perform better at the human-like vocabulary stages of tokenization (see Figure 3), by increasing the accessibility of useful distributions, and potentially also by providing suitable receptacles as semantic primitives for information derived from human cognition. As noted above, BPE can surface morphemes, words, and even multi-word or phrase-like units under different conditions, hinting that the specific efficiency pressures sometimes associated with word formation may apply across broader linguistic scales (recall Figures 2627Zipf (1949, 1965); Mandelbrot (1953, 1954). [^35] Perhaps across morphemes, words, phrases, sentences, and even larger conversational units. Where not already done, formalizing and computationally modeling these multi-scale efficiency effects represents an intriguing direction for future work. Extending this view (of how the word came to be) to LLM performance, perhaps structures like the word represent a stable solution to an informational optimization problem, one that balances the cost of expression, expressive power, [^36] Another way to say this might be: balancing the cost of transmitting – through koinos kosmos from one idios kosmos to another – with fidelity to the message. and learnability for both humans and machines. [^37] At least, when they must learn from human language. We discuss an LLM implementation of language – which, for example, requires less or at least different redundancy, as LLMs are not plagued by noisy rooms – and language as an abstract technology more in Zimmerman (2025). And if something like a word is the unit of language best suited to our ability to learn distributionally from the linguistic signal, then LLMs should perform best when provided easy access to plentiful word tokens.
@@ 2.3 Extispicy: exemplar tokens
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^57] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions585858e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Figure 41).
@@ 3.3 Flattening
The tokens form the semantic primitives for the model. They likely need to be well-formed receptacles for the kind of semantic information that makes up human world models, as that is a primary role of words for us.848484“A primary function of the lexicon is to encode novel concepts and ideas as they emerge over time” Xu et al. (2024). The Sapir-Whorf hypothesis in general, nameability (Lupyan’s Label-feedback hypothesis), and even theories from language acquisition like fast-mapping offer explanations for what kind of processes might need such structures Scholz et al. (2024); Lupyan (2012); Gelman and Brandone (2010). [^85] Although LLMs have something like MVP language, how that maps to the intersubjective or intrasubjective competence humans derive from language is unknown.
@@ 3.5 Bias, alignment, and ethics
Given the previous points, it is worth flagging that although they make up small portions of the total vocabulary, it is not uncommon for bad words to appear as tokens (especially since many new models remix or fine-tune existing models, and often incorporate an existing vocabulary off-the-shelf, without close inspection). [^98] LLMs are “so expensive to retrain from scratch that they are typically initialized with pre-trained models” Shumailov et al. (2024). In addition, changing the vocabulary would typically mean throwing away the initial training that had been done Bostrom and Durrett (2020). For example, 126 (out of 965) “bad words” are tokens in GPT-4o. We found some tokens that seem particularly likely to be vehicles for information that would work counter to alignment goals; for example, “cunt”, “slut”, “bitch”, “negro”, and “coon” are GPT-4o tokens, as are “iggers”, “igger”, and “IGGER”. The last three are representative of a class of tokens under suspicion as components of unsavory words, which, as aforementioned, could still carry that same content. [^99] In more detail: consider “igger”, “iggers”, and “IGGER” in GPT-4’s vocabulary (cl100k_base). Of course, those can be subwords of multiple words. For example, “bigger” and many variations of “trigger” are tokens, while “digger” is not. If “igger”, etc., is a useful subword frequently seen in other words, it would be included as a token. Depending on the details of the tokenization algorithm, inclusion of words containing it may or may not continue to contribute to its presence in the final vocabulary. That is, “igger” would be added to the vocabulary at some point, then “trigger” and “bigger”, and then, depending on the algorithm, a process may or may not remove “igger” if its value has decreased. However, there are a relatively small number of words containing it as a subword—some of those words are variants of the n-word, and several of the other, more mundane words are present as their own tokens. The presence of the subword “IGGER” in an entirely capitalized form is particularly suspicious (of low-quality or problematic content100100100Potentially up to and including the repulsive dregs of our collective written works - racism, hatred, bigotry, etc. related to the n-word or to “triggered”) given orthographic conventions on the internet.101101101Meaning both emphatic capitalization and the use of capitalization in memes. (See Table 11, Figures 3738.) The suggested content here is compatible, in our view, with features found in Templeton, et al., 2024: their analysis of the Claude 3 Sonnet transformer model at the feature level—above the token level—turned up concerning categories we could imagine being amongst just the sort of data we’re worried about making its way from the internet into the model, including “Phrases indicating profanity, vulgarity, obscenity or offensive language”; “Racist slurs and offensive language targeting ethnic/racial groups, particularly the N-word”; “Derogatory slurs, especially those targeting sexual orientation and gender identity”  Templeton et al. (2024).102102102And “Offensive, insulting or derogatory language, especially against minority groups and religions”; “Racist claims about crime”; “Mentions of violence, malice, extremism, hatred, threats, and explicit negative acts”. They also found information on famous people (which goes to a discussion of proper nouns in this paper and celebrities and Twitter in the in-progress Cramer, et al. Cramer et al. (2025)). They report that the features they found respond to and cause abstract behaviors Templeton et al. (2024). We think it is plausible that these features are, at least (and probably at most) in part, due to the mechanisms we’ve described.
@@
These same issues (bias, harmful content) can reasonably be assumed to persist throughout the model, since they are also attested in output, even after alignment. Hofmann et al., 2024 found that “HF [human feedback] training obscures the racism on the surface, but the racial stereotypes remain unaffected on a deeper level” Hofmann, Kalluri, and Jurafsky (2024). It is worth noting that many current alignment strategies would not be expected to fully redress the pathways for unwanted content that we’ve described.103103103Current prominent alignment strategies such as RLHF (Reinforcement Learning from Human Feedback) are unlikely to address the initial embedding matrix, although they can impact the content of the token vectors as they move through the model, since they are aimed at adjusting the model’s weights Stiennon et al. (2022). Even assuming alignment went exactly as intended and was as successful as possible, unwanted information could plausibly find a backdoor into the model through the initial token vectors during early training/ tokenization and remain untouched. A similar argument is put forward in Lee et al., 2025 with respect to the inadequacy of current unlearning techniques: “Even when achieving perfect behavioral agreement with an unlearning oracle, models retain latent capabilities that can be easily restored through finetuning” Lee et al. (2025).
@@
Beyond alignment, based on the the in-domain problem and the ouroboros: archaic, obscure, and obsolete words are likely not going to be tokenized.104104104Our results are consistent with this finding from Zhemchuzhina et al., 2022 Zhemchuzhina, Filippov, and Yamshchikov (2022)! How is that going to contribute to loss of beautiful cultural artifacts? Beyond buildings and even texts, language itself is a heritage site.105105105For example, peruse the Phrontistery Chrisomalis (1996-2021)! Similar arguments have been made at the narrative and cultural level, beyond the individual word or phrase: Generative AI acts as a “crucible of mediocrity” Ruiz, Zimmerman, and team (2023), as a regression to the mean Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).106106106We speculate that the probability mass of the logits corresponding to the subset of tokens involved in inference is likely to be relatively high, and that this is related to why models sometimes behaviourally collapse into inane repetition. More broadly, that repetition is a symptom of the iterative concentration of future plausibility based on previous plausibility, which is effectively a one-way ratchet Zimmerman et al. (2024). Imagine a hypothetical LLM generating output based on every previous utterance, including what it, itself, has said. As an analogy, compare sequence A005150 Foundation Inc. (2024a) to sequence A363054 Foundation Inc. (2024b): the initial seeds are obviously different, but as the output length grows, those differences are essentially lost to time (history being deeper on the right side of each term), and the sequences feel increasingly indistinguishable, locked in to reiterating the same sorts of patterns. It tends towards being normative and generic, which is undercut by creativity, novelty, and diversity Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).107107107“Beyond achieving linguistic coherence, beyond resolving the ‘hallucinations’ and errors, it is the performance of the generic and the normative, sounding ‘right,’ that sells AI tools… As currently designed, generative AI tend towards being generic, centrist, normative, and banal. Hallucinations undermine that, but so do cultural and narrative diversity” Gillespie (2024). This is, of course, partly a consequence of the fuzzing up of grammaticality or frequency with truth Zimmerman et al. (2024)!108108108McCoy et al., 2024, provides support as to (1) our characterization of plausible text generation as intertwining frequency and truth and (2) the theoretical importance of being able to modulate the primacy of the plausible text generation objective function, and the limitations of our current attempts in that vein NTP vs “reasoning”, efficacy of attempts to modulate primacy of NTP: “Specifically, o1 – like previous LLMs – is sensitive to the probability of examples and tasks, performing better and requiring fewer ‘thinking tokens’ in high-probability settings than in low-probability ones. These results show that optimizing a language model for reasoning can mitigate but might not fully overcome the language model’s probability sensitivity” McCoy et al. (2024b).,109109109For more discussion of the DH, and regime 1 vs. regime 2 learning (pattern-matching vs. learning from abstractions; hearing a novel word in context vs. reading a dictionary definition), see  Zimmerman (2025).
@@
Current LLMs have trouble with unusual data, at all scales.110110110Language, too, has this tendency inalienably within it, via its fungibility. Information technologies homogenize: that is both the benefit and price of fungibility. Fungibility is inherently degrounding, and degrounding is inherently homogenizing; it entails a loss of information or meaning, but a gain in utility. See first author’s thesis for more discussion of the construction of meaning along the fungibility-grounding spectrum Zimmerman (2025). Unusual data, data in the tails of the distribution, is important, both for cultural reasons and practical reasons of model performance and collapse. This is especially worth considering in the light of proposals to use LLMs to generate synthetic data (the ouroborous).111111111Shumailov et al., 2024 found that “in learning tasks in which the tails of the underlying distribution matter, one needs access to real human-produced data” as opposed to relying on LLM-generated data. They found that “access to the original data distribution is crucial” in order to prevent model collapse in subsequent generations Shumailov et al. (2024). Then, an initial tokenization step that curtails a model’s access to the tails of the attested word use distribution by omitting longer and rare words could also contribute to model collapse or other performance issues. Tokenization is a deterministic (though context-dependent) process, from within the perspective of a single algorithm being applied to a single fixed corpus. But tokenization as a cultural artifact in a larger system is less transparent.112112112Research continues to emerge that provides additional support for this argument Yakura et al. (2025).